emmagrace231 (emmagrace231):

Can someone help me with this? I don't understand this or how to do it. I have an exam after this lesson that is due by midnight! Will medal and fan for quick accurate responses!

OpenStudy (anonymous):

what is your question??

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

1. Is elimination

OpenStudy (anonymous):

oh i am sorry i did not relise what this was :( im only 6th grade

OpenStudy (anonymous):

2. Is A

OpenStudy (anonymous):

(2,1)

emmagrace231 (emmagrace231):

but how do you figure it out? I have to take an exam in a few min and this will be on it

OpenStudy (anonymous):

you see a \(-7x\) and a \(+7x\) so when you add them you will get no x term that is why elimination is easiest

emmagrace231 (emmagrace231):

but how do you figure out the second?

OpenStudy (danjs):

you can add equations together and get a new one here it will reduce to 1 variable, from what satellite said

OpenStudy (danjs):

that doesnt change anything for the solutions

emmagrace231 (emmagrace231):

i still don't understand. Can someone walk me through it?

OpenStudy (danjs):

you can solve it any way you want to, it is set up to have you prpbably do the elimination way. when you add the two equations, one variable will go to zero, you now have 1 equation and 1 variable. you can solve that for the variable. then use that to get the other

OpenStudy (mathmale):

Please note: Just giving away what you think is the answer is NOT allowed and is not helpful anyway.

OpenStudy (danjs):

need to at least try out a few before learning during the test.

OpenStudy (danjs):

a*x + b*y = c d*x + e*y = f ---------------adding (a+d)*x + (b + e)*y = c + f the point to have a+d or b+e be a zero, and removes that variable

OpenStudy (mathmale):

The two given equations are: 7x + 5y = 19 -7x - 2y =-16 -------------- Emma: Please add these two equations together right now. Think: What is 7x-7x? What is 5y-2y? What is 19-16?

OpenStudy (mathmale):

Note, Emma, that DanJS is correct when he writes, "when you add the two equations, one variable will go to zero, you now have 1 equation and 1 variable. you can solve that for the variable. then use that to get the other"

emmagrace231 (emmagrace231):

Can anyone just please help walk me through it? Bc I get that you are trying to help but I don't understand what you mean.

OpenStudy (mathmale):

I'm certain that you can combine 7 and -7. What does that give you? combining 7x and -7x is scarcely any harder. Emma: 7 -7 ------ ?

OpenStudy (danjs):

the lines are in a form of some number of x + some number of y = a number A*x+B*y = C A,B,C are just constants, numbers

OpenStudy (danjs):

if you add two of those forms of equations together, you can combine the like terms total x + total y = total constant number

OpenStudy (mathmale):

Emma: Examples: 2-2=0 139-139=0 -238 + 238 = 0 7 - 7 = ??

emmagrace231 (emmagrace231):

0

OpenStudy (danjs):

like i showed a bit back, here you see the total x's you get are a 7x and a -7x, totals to nothing

OpenStudy (mathmale):

Emma: see what Dan has just typed in. 7x - 7x = ???

emmagrace231 (emmagrace231):

7x-7x=0

OpenStudy (mathmale):

OK. that's right! So all you have left is: 5y = 19 -2y=-16 ----------- Combine the expressions in the left column. Your answer?

emmagrace231 (emmagrace231):

5y-2y=3y Do i Need to combine the 16 and -19 too?

emmagrace231 (emmagrace231):

@mathmale

emmagrace231 (emmagrace231):

@ganeshie8 @nincompoop @Hero @TheSmartOne

emmagrace231 (emmagrace231):

Please answer

ganeshie8 (ganeshie8):

Yes, your goal is to isolate \(y\)

ganeshie8 (ganeshie8):

5y = 19 -2y=-16 ----------- 3y = 3

OpenStudy (mathmale):

Now, Emma, divide both sides by 3. What final equation do you get?

ganeshie8 (ganeshie8):

\(3y = 3\) do you know how to isolate \(y\) in above equation ?

emmagrace231 (emmagrace231):

Don't you divide each side by the 3?

ganeshie8 (ganeshie8):

Do it and see...

emmagrace231 (emmagrace231):

y=1

emmagrace231 (emmagrace231):

right?

ganeshie8 (ganeshie8):

Perfect! you may look at options now

emmagrace231 (emmagrace231):

@ganeshie8

emmagrace231 (emmagrace231):

How do i get the answer 2,1 out of y=1?

OpenStudy (mathmale):

I'll let you answer that. You should write (2,1) and not just 2,1. What does (2,1) represent? Think.

OpenStudy (mathmale):

Separately, what does that '1' represent? Note that y ou solved those two equations for y and you got the result y=1.

emmagrace231 (emmagrace231):

does x=2 and y=1? That is the only explanation i can think of

OpenStudy (mathmale):

Yes, that is correct, altho I'd type, "y=1 when x=2" Very good. Now, we need to solve for x. Go back to the original 2 equations. Pick one of them. Type the whole equation in here.